OpenStudy (anonymous):

Two equal circles, O and Q of radius 8 intersect at X and Y. If OQ is 12, find the length of the common chord XY.

OpenStudy (cwrw238):



OpenStudy (cwrw238):

let R be the intersection of OQ and XY R is midpoint of XY XQ = 8, RQ = 1/2 * 12 = 6 use pythagoras to Find length of XR XY = 2 XR

OpenStudy (cwrw238):

do you follow kmeds?

OpenStudy (anonymous):

use right triangle properties.

OpenStudy (anonymous):

XY and OQ are perpendicular.

OpenStudy (anonymous):

@cwrw238 \[8\sqrt{3}\] ? O_-

OpenStudy (anonymous):

No...it's close...

OpenStudy (anonymous):



OpenStudy (anonymous):

oh no, hahaha. 64-36 lol

OpenStudy (cwrw238):

RQ^2 = 8^2 - 6^2 = 64-36 = 28

OpenStudy (anonymous):

4\sqrt{7}?\]

OpenStudy (anonymous):

\[4\sqrt{7}\] ?

OpenStudy (anonymous):

\[ XY=4\sqrt{7}\] if you got it correctly.

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

i was thinking 4. lol so i did, 64-16. hahahaha xD thank you @NewbieCarrot & @cwrw238 !